You can afford a $250 per month car payment. You've found a 5 year loan at 4% interest. How big of a loan can you afford?
Answer #1

Ans $13574.77

P = Regular Payments
PV = Loan Amount
r = rate of interest
n = no of periods
P = r (PV)
1 - (1 + r )-n
250 = (4.00%/12)*PV
1 - (1 / (1 + 4.00%/12)^60))
250 = 0.00333333333333333 * PV
0.180996896
PV = 250 * 0.180996896337451/ 0.00333333333333333
PV = 13574.77